OpenStudy (meowlover17):

Celina had a rectangular tablecloth with dimensions x feet and 3 feet. She put lace along 4 feet of the edge of the tablecloth. An expression is shown below: (2x + 6) - 4 Part A: What does the expression represent? (4 points) Part B: What does 2x + 6 represent? (3 points) Part C: What does 2x represent? (3 points)

OpenStudy (z4k4r1y4):

if you caculating the perimeter of the cloth what would you write?

OpenStudy (meowlover17):

Im not sure.

OpenStudy (z4k4r1y4):

 you would write: x + x +3 + 3

OpenStudy (meowlover17):

Yes, so (2x+6) represents the entire perimeter?

OpenStudy (z4k4r1y4):

yes

OpenStudy (meowlover17):

So does (2x+6)-4 represent the area or...?

OpenStudy (z4k4r1y4):

no. (2x + 6) is the entire perimeter...but 4 ft have lace on them.

OpenStudy (z4k4r1y4):

so what would (2x +6) - 4 be?

OpenStudy (meowlover17):

I'm not sure, but i'd have to guess, its the perimeter minus the lace. im not sure what it is otherwise.

OpenStudy (meowlover17):

Is there a more detailed way of describing the expression?

OpenStudy (z4k4r1y4):

you're right. it means that's the length of the perimeter that doesn't have lace.

OpenStudy (meowlover17):

Okay i see now, So for part A i'd put Part A: The expression (2x+6)-4 represents the part of the perimeter without the lace. Part B: (2x+6) Represents the entire perimeter, x+x being two sides, and 3+3 each being another two sides. Part C: 2x just represents the parallel sides of the rectangle. Also known as the sides opposite to each other, and would basically give its length once x is plugged in.

OpenStudy (z4k4r1y4):

couldn't have put it better myself!!

OpenStudy (meowlover17):

Thanks for your help, do you think you can help me with a few more?

OpenStudy (z4k4r1y4):

i can try

OpenStudy (meowlover17):

Okay here i go, Veronica teaches French in groups of four students. She rents a room for $200 per month. She charges $45 per hour per student. Veronica teaches seven groups of students, and her net earnings per month are $6,100. Part A: If each student learns French for the same number of hours, write an equation that can be used to calculate the number of hours each student learns French in a month. (5 points) Part B: How many hours did each student learn French each month? Show your work and justify each step of your work. (5 points)

OpenStudy (meowlover17):

@Z4K4R1Y4 you still there?

OpenStudy (z4k4r1y4):

yh

OpenStudy (z4k4r1y4):

Sorry, i can't help with this one.

OpenStudy (anonymous):

28*45*h=6100+200 28 represents the students 45 represents the cost per hour. h represents the # of hours

OpenStudy (meowlover17):

Its fine thanks for your help on the other one.

OpenStudy (z4k4r1y4):

np

OpenStudy (z4k4r1y4):

looks like you got help :)

OpenStudy (meowlover17):

@LegendarySadist What do i use that for?

OpenStudy (anonymous):

It is the answer for Part A and the formula you would use to solve Part B

OpenStudy (meowlover17):

Okay, so i just solve for h, in 28*48h=6,100+200?

OpenStudy (anonymous):

Honestly I'm not too sure about the "+200" part since I'm not very well acquainted with what "net earning" is referring to. You may or may not need the +200

OpenStudy (meowlover17):

i get 28*48h=6,300 --> 225 divded by 48 recieving 4.6875

OpenStudy (meowlover17):

Is that the answer for part b?

OpenStudy (anonymous):

ok, I confirmed the +200, so yeah that should be the answer.

OpenStudy (meowlover17):

but wait would the equation look like this (7)(4)(48)x+200=6,100?

OpenStudy (meowlover17):

instead of the +200 being on the other side

OpenStudy (anonymous):

Wait, wait

OpenStudy (meowlover17):

oh nevermind i think you are right

OpenStudy (anonymous):

I just noticed you made a typo and put 28*48 instead of 28*45

OpenStudy (meowlover17):

oops

OpenStudy (meowlover17):

so lemme try this again

OpenStudy (meowlover17):

28*45*h=6100+200 28 represents the students 45 represents the cost per hour. h represents the # of hours 1260h=6,300 h=5

OpenStudy (meowlover17):

Thanks i believe that is correct!

OpenStudy (anonymous):

Yeah, that looks right to me. Glad i caught that =)

OpenStudy (meowlover17):

Yea thanks, do you you think you have time for a few more?

OpenStudy (anonymous):

I could help with one or two, but no more than that

OpenStudy (meowlover17):

Okay sounds good

OpenStudy (anonymous):

Although it mostly depends on how quickly we can get through them

OpenStudy (meowlover17):

Sam is observing the velocity of a car at different times. After two hours, the velocity of the car is 54 km/h. After four hours, the velocity of the car is 58 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. (5 points) Part B: How can you graph the equation obtained in Part A for the first six hours? (5 points)

OpenStudy (meowlover17):

They should be rather fast

OpenStudy (meowlover17):

That is the next question

OpenStudy (anonymous):

Sorry for the delay, but i suppose it would be y=2x+50 wouldn't it? Or just 2x+50

OpenStudy (meowlover17):

Im not so sure.

OpenStudy (anonymous):

Well it satisfies both of them. For x of 2, you would get 54 and an x of 4 would get you 58

OpenStudy (meowlover17):

Okay but isnt that in slope intercept form

OpenStudy (meowlover17):

So X represents the amount of Hours, and Y represents different times???

OpenStudy (anonymous):

Slope intercept form is y=mx+b with m being the slope and b being the y-intercept. We have y=2x+52 with 2 being the m and 52 being the b

OpenStudy (meowlover17):

How can we put that in standard form??

OpenStudy (anonymous):

Oh, didn't notice that =(

OpenStudy (meowlover17):

does it being in slope intercept form change our entire equation? or can we just switch it to standard? also what do the X and Y variables represent

OpenStudy (anonymous):

I'd think it would just be -2x+y=50

OpenStudy (anonymous):

But honestly I'm not sure

OpenStudy (meowlover17):

Okay and x represents the velocity while y represents time? in other words we could do -2v+t=50???

OpenStudy (anonymous):

Standard form is just so useless that i never use it and don't remember how to convert.

OpenStudy (meowlover17):

Okay well how would i answer for part A: and B:????

OpenStudy (anonymous):

Actually yeah, just graphed them and they both fit so -2v+t=50 would be it

OpenStudy (meowlover17):

For part A do i put Part A: -2v+t=50 If you input 2 for velocity or 4 both satisfy what is given in the primary equation. v represents the velocity, while t represents time. To solve for v just add it to the other side --> t=+2(2)+50 --> t=54, then you could switch out 4 for v receiving 58. Part B: To graph, just plug in 0 for t and then 6 for t...you get (0,50) and (6,62) Connect those 2 points and you have your graph for the first 6 hours

OpenStudy (meowlover17):

So would that be correct? or is my part a or b wrong.

OpenStudy (anonymous):

Yeah, I'd say that looks good

OpenStudy (meowlover17):

Okay thanks do you think you have time for one more??

OpenStudy (anonymous):

Yeah, I can do one more

OpenStudy (anonymous):

Gonna be the last tho

OpenStudy (meowlover17):

Okay sounds great

OpenStudy (meowlover17):

here i go

OpenStudy (meowlover17):

The table below shows the outputs y for different inputs x: Input (x) 1 1 3 3 Output (y) 4 2 12 6 Part A: Does the table represent y as a function of x? Justify your answer. (5 points) Part B: The total cost f(x), in dollars, for renting a paddleboat for x hours is shown below: f(x) = 20 + 10x What is the value of f(120), and what does f(120) represent? (5 points)

OpenStudy (meowlover17):

I believe for Part a it would be Part A: No the GRAPH does NOT represent a function being there are more than one of the same x variable. But im not sure for part B

OpenStudy (meowlover17):

Sorry i meant the table not graph

OpenStudy (anonymous):

Wait, the input of 1 has two outputs? Definitely not a function there.